OpenStudy (petlovers123):

How many ounces of trail mix are in a bag that weighs 0.681 kilograms? Input only numeric values. (1 pound = 0.454 kg and 1 pound = 16 ounces) ____________________ @ganeshie8 @zepdrix

OpenStudy (petlovers123):

thank you for coming i need like major help

ganeshie8 (ganeshie8):

so you want to convert 0.681 kilograms to ounces is it ?

OpenStudy (petlovers123):

yea i guess so

ganeshie8 (ganeshie8):

you're given that 1 pound = 0.454 kg

ganeshie8 (ganeshie8):

\[0.681\,\color{red}{kg}\times \dfrac{1\,\color{blue}{pound}}{0.454\,\color{red}{kg}}\]

ganeshie8 (ganeshie8):

you may think of it like this : cancel kg top and bottom and the units change to pounds

ganeshie8 (ganeshie8):

similalry, use the other conversion factor to change pounds to ounces

OpenStudy (petlovers123):

how do i do that? im only in 6 grd

ganeshie8 (ganeshie8):

\[0.681\,\color{red}{kg}\times \dfrac{1\,\color{blue}{pound}}{0.454\,\color{red}{kg}}\times \dfrac{16\,\color{purple}{ounces}}{1\,\color{blue}{pound}}\]

ganeshie8 (ganeshie8):

cancel the kg in red cancel the pound in blue you will be left with ounces

ganeshie8 (ganeshie8):

\[\require{cancel} 0.681\,\color{red}{\cancel{kg}}\times \dfrac{1\,\color{blue}{\cancel{pound}}}{0.454\,\color{red}{\cancel{kg}}}\times \dfrac{16\,\color{purple}{ounces}}{1\,\color{blue}{\cancel{pound}}}\] \[0.681\times \dfrac{1}{0.454}\times \dfrac{16\,\color{purple}{ounces}}{1}\]

ganeshie8 (ganeshie8):

simplify

ganeshie8 (ganeshie8):

\[ \dfrac{0.681\times 16}{0.454}\, \color{purple}{ounces}\]

OpenStudy (petlovers123):

ok